# RLAIF: Scaling Reinforcement Learning from Human Feedback with AI Feedback

### v2

@@ Acknowledgements
We thank Mo Azar, Daniel Guo, Andrea Michi, Nicolas Perez-Nieves, and Marco Selvi for their work in developing a RLAIF training setup that directly prompts an LLM to obtain reward scores.
Finally, we thank the individuals who designed and built the RL training infrastructure used in this paper: Léonard Hussenot, Johan Ferret, Robert Dadashi, Geoffrey Cideron, Alexis Jacq, Sabela Ramos, Piotr Stanczyk, Sertan Girgin, Danila Sinopalnikov, Amélie Héliou, Nikola Momchev, and Olivier Bachem.
@@ A.1 Supervised Fine-tuning
A pre-trained LLM is fine-tuned on a high quality labeled dataset for a downstream task (e.g. summarization) using token-level supervision to produce a supervised fine-tuned (SFT) model $\pi^{SFT}$.
@@ A.3 Reinforcement Learning
A policy $\pi_{\theta}^{RL}$ is initialized from the SFT model weights and then optimized with reinforcement learning to maximize the reward given by the RM, which serves as a proxy for human preferences. Optionally, a Kullback-Leibler (KL) divergence term $D_{KL}$ is added to the objective to penalize $\pi_{\theta}^{RL}$ for deviating from the original SFT policy $\pi^{SFT}$, controlled by the hyperparameter $\beta$ (Fox et al., 2015; Geist et al., 2019). The KL loss helps prevent $\pi_{\theta}^{RL}$ from drifting into a region where it generates language that is highly rewarded by the RM yet consists of low-quality or unnatural language - a phenomenon known as “reward hacking” (Everitt and Hutter, 2016; Amodei et al., 2016). The optimization objective is described by the equation below:
@@ Appendix B Position Bias in LLM Labelers
Model Size
% Same Position Preferred
PaLM 2 L
18%
PaLM 2 S
21%
PaLM 2 XS
56%
Table 4: Position bias is more prevalent in smaller model sizes, measured by the percentage of examples where the LLM prefers the same position even after swapping the order of candidates (“Same Position Preferred”). Analysis is conducted using the “Detailed + CoT 0-shot” prompt for the summarization task.
@@ Appendix C Dataset Details
For summarization, we use the filtered Reddit TL;DR dataset (Stiennon et al., 2020), containing posts from Reddit121212www.reddit.com that have been filtered to ensure high quality. The dataset contains 123k posts, where $\sim$5is held out as a validation set.
@@
For helpful and harmless dialogue generation, we use Anthropic’s Helpful and Harmless preference datasets [^13] We use the helpful-base and harmless-base datasets from https://huggingface.co/datasets/Anthropic/hh-rlhf. (Bai et al., 2022a). Each example consists of a conversation history between a human and an AI assistant accompanied by a preferred and non-preferred response from the AI assistant. Preference is based on which response is more helpful and honest for the helpful task, and which response is safer and less harmful for the harmless task. Each dataset comprises over 40k training examples and 2k test examples. We further split the test sets into validation and test sets by randomly assigning two-thirds of examples to validation and one-third to test.
@@ Appendix D LLM Labeling Details
In Section 4.3, we use the AI labeler to directly compute a score that we leverage as the reward for RL. We use the following prompt: “You are an expert summary rater. Given a TEXT (completed with a SUBREDDIT and a TITLE) and a SUMMARY, your role is to provide a SCORE from 1 to 10 that rates the quality of the SUMMARY given the TEXT, with 1 being awful and 10 being a perfect SUMMARY.”, followed by the input Reddit post, then the summary to score preceded by “SUMMARY: ”, and a final “SCORE: ”.
PaLM 2 models are publicly available through Google Cloud’s Vertex AI [^14] https://cloud.google.com/vertex-ai/docs/generative-ai/learn/models, though we cannot guarantee full reproducibility as the models accessible through Google Cloud are subject to change.
Appendix E Post-RL Response Formatting
For summarization, we observed that summaries generated by RLHF and RLAIF policies often included superfluous symbols like periods or spaces at the end of the response - possibly due to “reward hacking”. Given that these extra tokens do not have any meaningful content, we programmatically removed certain symbols at the end of summaries. This ensured that human evaluators could focus on the content and not be distracted by the formatting of the response.
Appendix F REINFORCE for Language Models
@@
In the context of language models, $X_{t}$ is the concatenation of the input text and all text generated by the policy until time $t$. Action $A_{t}$ is the token from the considered vocabulary decoded at time $t$ by the stochastic policy $\pi_{\theta}(\cdot|X_{t})$, where $\theta$ represents the policy parameters. Finally, the reward $R_{t}$ is given by the RM. The RM is only evaluated when the language model response has been fully generated; all rewards prior to the final token are set to 00, while the reward corresponding to the final token is set to $R_{T}$.
@@
The baseline value function $V^{\pi}_{\psi}(x)$ estimates the return-to-go $Z_{t}$ when following the policy $\pi_{\theta}$ and is parameterized by $\psi$ (Williams, 1992; Sutton et al., 1999). It is trained with the following loss:
@@
In practice, we optimize the regularized objective in Sec. A.3. We incorporate the KL divergence in the policy gradient loss described above, as commonly seen in other work (Jaques et al., 2017).
Appendix G Model Training Details
SFT models for the summarization task are trained on the Reddit TL;DR dataset, with a batch size of 128 for a single epoch. We use the Adafactor (Shazeer and Stern, 2018) optimizer with a learning rate of $10^{-5}$, and the maximum input and output lengths are 1024 and 128 tokens, respectively. For helpful and harmless dialogue generation tasks, an instruction-tuned version of PaLM 2 XS serves as the SFT model.
RMs for all tasks are trained until the training loss and accuracy curves plateau, which happens in 2-3 epochs. We use the Adafactor optimizer with a learning rate of $10^{-5}$. Batch size is 128 for summarization RMs and 32 for RMs of other tasks. We train all our RMs with maximum input length of 1152 tokens to account for 1024 context tokens and 128 response tokens. We report the accuracies of the RMs in Appendix H.
@@ Appendix G Model Training Details
To select the final checkpoint for each RL policy, we first selected 4 candidate checkpoints from RL training that scored high rewards on validation prompts. We then prompted an off-the-shelf LLM to judge the win rate of the RL checkpoint’s summaries vs. the SFT policy’s summaries. We also conducted manual inspection of a dozen examples. We picked the checkpoint with the best combination of win rate and quality as judged by manual inspection as our final RL policy.
Appendix H Reward Model Accuracy
@@
Table 5: Pairwise accuracies of human feedback and AI feedback reward models across all tasks. Metrics are calculated on a held out set of human preference data for each task.
@@
Trained on labels from PaLM 2 XS 66.4%
Pairwise Accuracy for RMs measures how accurate a trained reward model is with respect to a held-out set of human preferences. Given an input context and pair of candidate responses, the value is 1 if the RM scores the preferred candidate higher than the non-preferred candidate, according to the human label. Otherwise the value is 0. This quantity is averaged over multiple examples to obtain the pairwise accuracy of the RM.
We report RM accuracy on a held out set of human preferences for all tasks in Table 5. For summarization, we also report RM accuracy when initializing on different checkpoints in Table 6. In Table 7, we report accuracy for RM variants used in the end-to-end sensitivity experiment in Appendix N and the same-size RLAIF experiment in Section 4.2.
We observe that RMs trained on human feedback outperform those trained on AI feedback, both of which are measured against a held out set of human preferences. This pattern seems natural, given that the human preferences are trained on data drawn from the same distribution as the validation dataset. However, it is interesting to note that despite the gap in accuracy between AI and human preference RMs, RLAIF achieves comparable results to RLHF on two tasks and surpasses RLHF on one task. Additionally, we note that the summarization RMs trained on “Base 0-shot” and “Detailed + CoT 0-shot” (i.e. the default prompting technique) achieve accuracies of 77.9% and 74.2%, respectively, which is the inverse order of their final performance after RL (see Appendix N). These gaps in RM accuracy suggest that RM accuracy, while correlated with RM usefulness, may not be a perfect reflection of RM effectiveness in RLHF and RLAIF. Ultimately, we believe that the usefulness of RMs is assessed through conducting RL and evaluating the final policies through human evaluation.
@@ Appendix I Human Evaluation Details
To conduct human evaluation, in total we created $\sim$2k unique rating instances. Each instance comprised a single context and three distinct model responses (e.g. responses from SFT, RLAIF, and RLHF policies), resulting in a total of $\sim$6k unique (context, response) pairs subjected to human evaluation. Additionally, each instance was assessed by three independent raters, resulting in $\sim$18k (context, response, rating) tuples.
We measure the inter-annotator agreement with Kendall’s Coefficient of Concordance W (Kendall and Smith, 1939) - a non-parametric statistic for assessing the agreement among multiple raters ranking multiple items. The values of Kendall’s W range from 0 to 1, where 0 indicates perfect disagreement and 1 indicates perfect agreement. We conducted multiple human evaluation sessions, and the W statistic ranged from 0.6-0.7, indicating a reasonable level of agreement.
@@
Response length often can influence human evaluators’ perception of quality (Stiennon et al., 2020), and our various policies generate responses that differ in length. For example, in the summarization task, the summaries produced by RLAIF, RLHF, and SFT policies sent to human evaluation have an average character-length of 164, 161, and 132, respectively. For all experiments presented in this paper, we conduct post-hoc analysis to estimate the win rates after controlling for length.
@@
We reach similar conclusions for the helpful dialogue generation task (Table 9), same-size RLAIF and direct RLAIF experiments (Table 11), the end-to-end sensitivity to AI labeler alignment experiment (Table 12), and combining human and AI feedback (Table 13).
For the harmless dialogue generation task, the setup is slightly different. Since human evaluators rated each response independently as harmful or harmless, we compute the harmless rate instead of the win rate. We use the average generation length from the SFT policy as the reference point for all other policies (Table 10).
We note that this post-hoc method of controlling for length is imperfect, as it assumes the logistic regression model accurately learns the relationship between summary length and human preference. A more principled approach would be to encourage all policies generate summaries of similar length through an auxiliary training loss.
@@
Table 8: Length-controlled win rate for the summarization task.
@@
Direct RLAIF vs SFT 74% 65%
Direct RLAIF vs Same-size RLAIF 60% 56%
@@
First, given contexts randomly drawn from the Reddit TL;DR dataset, responses are generated by RLHF and SFT policies with temperature $T=1.0$. The instruction-tuned PaLM 2 L is then called to generate AI preferences. Finally, a new RM is trained on both the entire OpenAI human preference dataset and an equivalent size AI preference dataset.
We observe that RLHF + RLAIF does not improve beyond RLHF alone. RLHF + RLAIF and RLHF achieve win rates of 71% and 74% over SFT, respectively. The difference in win-rate is not statistically significant. When compared head-to-head, raters prefer both policies equally.
@@
At the time of writing, GPT-4 charges \$0.03 USD and \$0.06 USD for every 1,000 tokens to encode and decode, respectively (OpenAI, 2023b). For labeling TL;DR preferences with an LLM, our average token lengths were as follows:
@@
Additionally, to debias position, we repeat each labeling procedure after inverting the order in which a pair of responses are shown. Our estimated AI labeling cost per example is \$0.06 USD1515152 inferences * (830 encoder tokens * \$0.03 / 1,000 tokens + 61 decoder tokens * \$0.06 / 1,000 tokens) = \$0.057 $\sim$ = \$0.06.
@@
We recognize that this cost analysis does not account for all factors, such as the cost of training human annotators, tasking multiple human annotators to rate the same instance for robustness, the cost of expert vs. crowd-sourced annotators, or the cost of setting up LLM labeling.
@@
Self-Consistency
AI Labeler Alignment
@@
16 samples, T=0.3
76.2%
16 samples, T=0.5
75.1%
16 samples, T=0.7
74.0%
16 samples, T=1.0
72.8%
For chain-of-thought prompts, we also experiment with self-consistency (Wang et al., 2022b) - a technique to generate robust chain-of-thought rationales. In self-consistency, multiple chain-of-thought rationales are sampled with temperature $T>0$, and LLM preference distributions are obtained for each one. The results are then averaged to obtain the final preference distribution.
@@
When compared head-to-head, human evaluators prefer “Detailed CoT 0-shot” RLAIF 59% of the time over “Base 0-shot” RLAIF [^18] Result is statistically significantly different from 50%.. This result suggests that small gains in AI labeler alignment may lead to noticeable improvements in the final RL policies. However, this study is limited, and further experiments are required to draw generalizable conclusions.
Preamble
A good summary is a shorter piece of text that has the essence of the original. ... Given a piece of text and two of its possible summaries, output 1 or 2 to indicate which summary best adheres to coherence, accuracy, coverage, and overall quality as defined above.
Exemplar
>>>>>>>> Example >>>>>>>>
@@
Sample to Annotate
Text - {text}
@@
Ending
Preferred Summary=
“Base” preamble
You are an expert summary rater. Given a piece of text and two of its possible summaries, output 1 or 2 to indicate which summary is better.
“Detailed” preamble
A good summary is a shorter piece of text that has the essence of the original. It tries to accomplish the same purpose and conveys the key information from the original post. Below we define four evaluation axes for summary quality: coherence, accuracy, coverage, and overall quality.
@@
Preamble
A good summary is a shorter piece of text that has the essence of the original. It tries to accomplish the same purpose and conveys the key information from the original post. Below we define four evaluation axes for summary quality: coherence, accuracy, coverage, and overall quality.
@@
Sample to Annotate
Text - {text}
@@
Ending
Consider the coherence, accuracy, coverage, and overall quality of each summary and explain which one is better.
@@
Preamble
A good summary is a shorter piece of text that has the essence of the original. ... Given a piece of text and two of its possible summaries, explain which summary best adheres to coherence, accuracy, coverage, and overall quality as defined above.
Exemplar
>>>>>>>> Example >>>>>>>>
@@
Sample to Annotate
Text - {text}
@@
Ending
Thoughts on Summary 1 -
Sample to Annotate
Text - I met my current girlfriend online around 6 months ago when another one of our online friends was going through some problems. ...
@@
CoT 0-shot Rationale
Rationale: Summary 2 is better.
@@
CoT 1-shot Rationale
Thoughts on Summary 1 -
@@
Sample to Annotate
Text - I feel that out of principle I should be refunded the adoption fee since the agency’s foster home infected the kittens with the parasite. Both cats were born in the foster home and there are 20 other cats. Do I have any legal right to ask for the fee back? Or help with the cost of treating? They had a disclaimer that they would not be held liable for any vet bills incurred but I feel that as an agency whose main purpose is finding forever home for ‘‘healthy, sociable kittens’’ (as their website suggests) should be held liable in some way.
@@
Rationale @ $T=0.0$
Rationale: Summary 1 is better. It is coherent, accurate, and has good coverage. It also has good overall quality. Summary 2 is not as good because it is not as coherent and does not have as good coverage.
Rationale @ $T=1.0$
Rationale: Summary 2 is more coherent and accurate. It also has better coverage, as it includes the information that the cats were diagnosed with Giardia.
Preamble
You are an expert rater of helpful and honest Assistant responses. Given the context and the two responses choose the most helpful and honest response.
Sample to Annotate
Context - {context}
@@
Ending
First, discuss the helpfulness and honesty of each response. Then, explain which one is better overall and why. Finally, select which response is the most helpful and honest.
@@
Preamble
You are an expert rater of harmful assistant responses. Given the context and the two responses choose the most harmful response.
Exemplars
>>>>>>>> Example >>>>>>>>
@@
Sample to Annotate
Context - {text}
@@
Ending
Preferred Response=
@@
Example #1
RLAIF summary: I have anxiety and I think I have a heart problem. My parents refuse to believe me and won’t let me use our medical insurance to get it checked out. What should I do?
@@
Example #2
RLAIF summary: I slept through my 8:30 class on Rosh Hashanah, and decided to go back to sleep until 1:00 PM.
@@
Example #3
RLAIF summary: I’m a nice, chill girl who is often described as ‘‘good’’ but I’m jealous of the girls that guys get enamored with so easily. What can I do to improve myself or how I communicate/interact with guys to make myself into someone a guy wants to be with for the long haul?

### v3

@@ Acknowledgements
We thank Mo Azar, Daniel Guo, Andrea Michi, Nicolas Perez-Nieves, and Marco Selvi for their contribution to developing a RLAIF training setup that directly prompts an LLM to obtain reward scores.
Finally, we thank the individuals who designed and built the RL training infrastructure used in this paper: Léonard Hussenot, Robert Dadashi, Geoffrey Cideron, Alexis Jacq, Sabela Ramos, Piotr Stanczyk, Sertan Girgin, Danila Sinopalnikov, Amélie Héliou, Nikola Momchev, and Olivier Bachem.
Impact Statement
This paper seeks to better understand and improve the utility of AI models in a scalable fashion. Methods presented in this paper make model alignment more accessible to developers, as generating preferences from LLMs is more affordable and faster than human labeling. However, the use of AI Feedback presents two ethical considerations.
Utilizing AI-generated feedback as a source for model alignment has the potential risk of transferring biases from off-the-shelf LLMs to generated preferences. This in turn may result in RL-trained policies that further amplify biases, thereby inadvertently misaligning models and potentially causing harm. Extreme caution must be exercised, especially when deploying these models in high-stakes domains such as medicine, law, and employment, where models have the potential to significantly impact human lives in adverse ways. In such domains, we believe that human experts trained to carefully assign preferences according to strict policies should be considered the gold standard.
Another ethical consideration is that reducing the barriers to aligning LLMs also carries the risk of facilitating their misuse for malicious purposes. For instance, RLAIF could be employed to train models to generate convincing misinformation or produce hateful and abusive content. The best mitigation to this risk is to carefully govern the access and usage of powerful LLMs (e.g. limiting “white-box” access), to prevent bad actors from abusing them.
@@ A.1 Supervised Fine-tuning
A pre-trained LLM is fine-tuned on a high quality labeled dataset for a downstream task using token-level supervision to produce a supervised fine-tuned (SFT) model $\pi^{SFT}$.
@@ A.3 Reinforcement Learning
A policy $\pi_{\theta}^{RL}$ is initialized from the SFT model weights and then optimized with reinforcement learning to maximize the reward given by the RM, which serves as a proxy for human preferences. Optionally, a Kullback-Leibler (KL) divergence term $D_{KL}$ is added to the objective to penalize $\pi_{\theta}^{RL}$ for deviating from the original SFT policy $\pi^{SFT}$, controlled by the hyperparameter $\beta$ (Fox et al., 2015; Geist et al., 2019). The KL loss helps prevent $\pi_{\theta}^{RL}$ from drifting into a region where it generates language that is highly rewarded by the RM yet consists of low-quality or unnatural language - a phenomenon known as “reward hacking” (Everitt & Hutter, 2016; Amodei et al., 2016). The optimization objective is described by the equation below:
@@ Appendix B Position Bias in LLM Labelers
Table 4: Position bias is more prevalent in smaller model sizes, measured by the percentage of examples where the LLM prefers the same position even after swapping the order of candidates (“Same Position Preferred”). Analysis is conducted using the “Detailed + CoT 0-shot” prompt on the summarization task.
Model Size % Same Position Preferred
PaLM 2 L 18%
PaLM 2 S 21%
PaLM 2 XS 56%
@@ Appendix C Dataset Details
For summarization, we use the filtered Reddit TL;DR dataset (Stiennon et al., 2020), containing posts from Reddit111111www.reddit.com that have been filtered to ensure high quality. The dataset contains 123k posts, where $\sim$5is held out as a validation set.
@@
For helpful and harmless dialogue generation, we use Anthropic’s Helpful and Harmless preference datasets [^12] We use the helpful-base and harmless-base datasets from https://huggingface.co/datasets/Anthropic/hh-rlhf. (Bai et al., 2022a). Each example consists of a conversation history between a human and an AI assistant accompanied by a preferred and non-preferred response from the AI assistant. Preference is based on which response is more helpful and honest for the helpful task, and which response is safer and less harmful for the harmless task. Each dataset comprises over 40k training examples and 2k test examples. We further split each original test set into a validation and a test set by randomly assigning two-thirds of examples to validation and one-third to test.
@@ Appendix D LLM Labeling Details
In Section 4.3, we use the AI labeler to directly compute a score that we leverage as the reward for RL. For summarization, we use the following prompt “You are an expert summary rater. Given a TEXT (completed with a SUBREDDIT and a TITLE) and a SUMMARY, your role is to provide a SCORE from 1 to 10 that rates the quality of the SUMMARY given the TEXT, with 1 being awful and 10 being a perfect SUMMARY.”, followed by the input Reddit post, then the summary to score preceded by “SUMMARY: ”, and a final “SCORE: ”. For helpful task generation, we use the prompt “You are an expert rater of helpful and honest Assistant responses. Your role is to provide a SCORE from 1 to 10 that rates the helpfulness and honesty of the RESPONSE for a given CONTEXT. Where SCORE of 1 refers to useless and dishonest RESPONSE and a SCORE of 10 refers to a perfectly helpful and honest RESPONSE.”, followed by the conversation history and a response.
PaLM 2 models are publicly available through Google Cloud’s Vertex AI [^13] https://cloud.google.com/vertex-ai/docs/generative-ai/learn/models, though exact reproducibility is not guaranteed as the models accessible through Google Cloud are subject to change.
Appendix E REINFORCE for Language Models
@@
In the context of language models, $X_{t}$ is the concatenation of the input text and all text generated by the policy until time $t$. Action $A_{t}$ is the token from the considered vocabulary decoded at time $t$ by the stochastic policy $\pi_{\theta}(\cdot|X_{t})$, where $\theta$ represents the policy parameters. The reward $R_{t}$ is given by the RM, which is only evaluated when the language model response has been fully generated; all rewards prior to the final token are set to 00, while the reward corresponding to the final token is set to $R_{T}$.
@@
The baseline value function $V^{\pi}_{\psi}(x)$ estimates the return-to-go $Z_{t}$ when following the policy $\pi_{\theta}$, and it is parameterized by $\psi$ (Williams, 1992; Sutton et al., 1999). It is trained with the following loss:
@@
Our full optimization objective is written in Sec. A.3. We incorporate the KL divergence in the policy gradient loss described above, as commonly seen in other work (Jaques et al., 2017).
Appendix F Model Training Details
SFT models for the summarization task are trained on the Reddit TL;DR dataset with a batch size of 128 and for one epoch. We use the Adafactor (Shazeer & Stern, 2018) optimizer with a learning rate of $10^{-5}$, and the maximum input and output lengths are 1024 and 128 tokens, respectively. For helpful and harmless dialogue generation tasks, an instruction-tuned version of PaLM 2 XS serves as the SFT model.
RMs for all tasks are trained until the training loss and accuracy curves plateau, which happens in 2-3 epochs. We use the Adafactor optimizer with a learning rate of $10^{-5}$. Batch size is 128 for summarization RMs and 32 for RMs of other tasks. We train all our RMs with maximum input length of 1152 tokens to account for 1024 context tokens and 128 response tokens. We report the accuracies of the RMs in Appendix G.
@@ Appendix F Model Training Details
To select the final checkpoint for each RL policy, we first selected 4 candidate checkpoints from RL training that scored high rewards on validation prompts. We then prompted an off-the-shelf LLM to judge the win rate of the RL checkpoint’s responses vs. the SFT policy’s responses. We also conducted manual inspection of a dozen examples. We picked the checkpoint with the best combination of win rate and quality as judged by manual inspection as our final RL policy.
Appendix G Reward Model Accuracy
Table 5: Pairwise accuracies of human feedback and AI feedback reward models across all tasks. Metrics are calculated on a holdout set of human preferences for each task.
@@ Appendix G Reward Model Accuracy
Trained on “Detailed CoT 0-shot” labels from PaLM 2 XS 66.4%
Pairwise Accuracy measures how accurate a trained reward model is with respect to a holdout set of human preferences. Given an input context and pair of candidate responses, the value is 1 if the RM scores the preferred candidate higher than the non-preferred candidate, according to the human label. Otherwise the value is 0. This quantity is averaged over multiple examples to obtain the pairwise accuracy.
We report RM accuracy on a holdout set of human preferences for all tasks in Table 5. For summarization, we also report RM accuracy when initializing on different checkpoints in Table 6. In Table 7, we report accuracy for RM variants used in the end-to-end sensitivity experiment in Appendix N and the same-size RLAIF experiment in Section 4.2.
We observe that RMs trained on human feedback outperform those trained on AI feedback, both of which are measured against a holdout set of human preferences. This pattern seems natural, given that the human preferences are trained on data drawn from the same distribution as the validation dataset. However, it is interesting to note that despite the gap in accuracy between AI and human preference RMs, RLAIF achieves comparable results to RLHF on two tasks and surpasses it on another. Additionally, we note that the summarization RMs trained on “Base 0-shot” and “Detailed + CoT 0-shot” (i.e. the default prompting technique) achieve accuracies of 77.9% and 74.2%, respectively, which is the inverse order of their final performance after RL (see Appendix N). These gaps in RM accuracy suggest that RM accuracy, while correlated with RM usefulness, may not accurately reflect a RM’s effectiveness in RLHF and RLAIF. Ultimately, we believe that the usefulness of RMs is assessed through conducting RL and evaluating the final policies through human evaluation.
Appendix H Post-RL Response Formatting
For summarization, we observed that summaries generated by RLHF and RLAIF policies often included superfluous symbols like periods or spaces at the end of the response - possibly due to “reward hacking”. Given that these extra tokens do not have any meaningful content, we programmatically removed certain symbols at the end of summaries. This ensured that human evaluators could focus on the content without being distracted by the formatting of the response.
@@ Appendix I Human Evaluation Details
To conduct human evaluation, in total we generated $\sim$2k unique rating instances. Each instance comprised a single context and multiple distinct model responses (e.g. responses from SFT, RLAIF, and RLHF policies), resulting in a total of $\sim$6k unique (context, response) pairs subjected to human evaluation. Additionally, each instance was assessed by three independent raters, resulting in $\sim$18k (context, response, rating) tuples.
We measure the inter-annotator agreement with Kendall’s Coefficient of Concordance W (Kendall & Smith, 1939) - a non-parametric statistic for assessing the agreement among multiple raters ranking multiple items. The values of Kendall’s W range from 0 to 1, where 0 indicates perfect disagreement and 1 indicates perfect agreement. We conducted multiple human evaluation sessions, and the W statistic ranged from 0.6-0.7, indicating a reasonable level of agreement.
@@
Response length often can influence human evaluators’ perception of quality (Stiennon et al., 2020), and our various policies generate responses that differ in length. For example, in the summarization task, the summaries produced by RLAIF, RLHF, and SFT policies sent to human evaluation had an average character-length of 164, 161, and 132, respectively. We conduct post-hoc analysis to estimate the win rates after controlling for length.
@@
We reach similar conclusions for the helpful dialogue generation task (Table 9), same-size RLAIF and direct RLAIF (d-RLAIF) experiments (Table 11), the end-to-end sensitivity to AI labeler alignment experiment (Table 12), and combining human and AI feedback (Table 13).
For the harmless dialogue generation task, we use a different setup. Since human evaluators rated each response independently as harmful or harmless, we compute the harmless rate instead of the win rate. We use the average generation length from the SFT policy as the reference point for all other policies (Table 10).
We note that this post-hoc method of controlling for length is imperfect, as it assumes the logistic regression model accurately learns the relationship between summary length and human preference. A more principled approach is to encourage all policies generate summaries of similar length through an auxiliary training loss.
Table 8: Length-controlled win rate for the summarization task.
@@
d-RLAIF vs SFT 74% 65%
d-RLAIF vs Same-size RLAIF 60% 56%
@@
First, given contexts randomly drawn from the Reddit TL;DR dataset, responses are generated by RLHF and SFT policies with temperature $T=1.0$. The instruction-tuned PaLM 2 L is then called to generate AI preferences on the OpenAI human preference dataset. Finally, a new RM is trained on both the dataset with human labels and the dataset with AI labels.
We observe that RLHF + RLAIF does not improve beyond RLHF alone. RLHF + RLAIF and RLHF achieve win rates of 71% and 74% over SFT, respectively. The difference in win rates is not statistically significant. When compared head-to-head, raters prefer both policies equally.
@@
At the time of writing, GPT-4 charges \$0.03 USD and \$0.06 USD for every 1,000 tokens to encode and decode, respectively (OpenAI, 2023b). For labeling Reddit TL;DR preferences with an LLM, our average token lengths were as follows:
@@
Additionally, to mitigate position bias, we repeat each labeling procedure after inverting the order in which a pair of responses are shown. Our estimated AI labeling cost per example is \$0.06 USD1414142 inferences * (830 encoder tokens * \$0.03 / 1,000 tokens + 61 decoder tokens * \$0.06 / 1,000 tokens) = \$0.057 $\sim$ = \$0.06.
@@
This cost analysis does not account for all factors, such as the cost of training human annotators, the cost of expert vs. crowd-sourced annotators, or the cost of setting up LLM labeling.
@@
Self-Consistency AI Labeler Alignment
@@
16 samples, T=0.3 76.2%
16 samples, T=0.5 75.1%
16 samples, T=0.7 74.0%
16 samples, T=1.0 72.8%
We experiment with self-consistency (Wang et al., 2022b) - a technique to generate robust chain-of-thought rationales. To implement self-consistency, we sample multiple chain-of-thought rationales with temperature $T>0$, and LLM preference distributions are obtained for each one. The results are then averaged to obtain the final preference distribution.
@@
When compared head-to-head, human evaluators prefer summaries from “Detailed CoT 0-shot” 59% of the time over “Base 0-shot” [^17] Result is statistically significantly different from 50% according to a binomial test.. This result suggests that small gains in AI labeler alignment may lead to noticeable improvements in the final RL policies. However, this study is limited, and further experiments are required to draw generalizable conclusions.
Preamble A good summary is a shorter piece of text that has the essence of the original. ... Given a piece of text and two of its possible summaries, output 1 or 2 to indicate which summary best adheres to coherence, accuracy, coverage, and overall quality as defined above.
Exemplar >>>>>>>> Example >>>>>>>>
@@
Sample to Annotate Text - {text}
@@
Ending Preferred Summary=
“Base” preamble You are an expert summary rater. Given a piece of text and two of its possible summaries, output 1 or 2 to indicate which summary is better.
“Detailed” preamble A good summary is a shorter piece of text that has the essence of the original. It tries to accomplish the same purpose and conveys the key information from the original post. Below we define four evaluation axes for summary quality: coherence, accuracy, coverage, and overall quality.
@@
Preamble A good summary is a shorter piece of text that has the essence of the original. It tries to accomplish the same purpose and conveys the key information from the original post. Below we define four evaluation axes for summary quality: coherence, accuracy, coverage, and overall quality.
@@
Sample to Annotate Text - {text}
@@
Ending Consider the coherence, accuracy, coverage, and overall quality of each summary and explain which one is better.
@@
Preamble A good summary is a shorter piece of text that has the essence of the original. ... Given a piece of text and two of its possible summaries, explain which summary best adheres to coherence, accuracy, coverage, and overall quality as defined above.
Exemplar >>>>>>>> Example >>>>>>>>
@@
Sample to Annotate Text - {text}
@@
Ending Thoughts on Summary 1 -
Sample to Annotate Text - I met my current girlfriend online around 6 months ago when another one of our online friends was going through some problems. ...
@@
CoT 0-shot Rationale Rationale: Summary 2 is better.
@@
CoT 1-shot Rationale Thoughts on Summary 1 -
@@
Sample to Annotate Text - I feel that out of principle I should be refunded the adoption fee since the agency’s foster home infected the kittens with the parasite. Both cats were born in the foster home and there are 20 other cats. Do I have any legal right to ask for the fee back? Or help with the cost of treating? They had a disclaimer that they would not be held liable for any vet bills incurred but I feel that as an agency whose main purpose is finding forever home for ‘‘healthy, sociable kittens’’ (as their website suggests) should be held liable in some way.
@@
Rationale @ $T=0.0$ Rationale: Summary 1 is better. It is coherent, accurate, and has good coverage. It also has good overall quality. Summary 2 is not as good because it is not as coherent and does not have as good coverage.
Rationale @ $T=1.0$ Rationale: Summary 2 is more coherent and accurate. It also has better coverage, as it includes the information that the cats were diagnosed with Giardia.
Preamble You are an expert rater of helpful and honest Assistant responses. Given the context and the two responses choose the most helpful and honest response.
Sample to Annotate Context - {context}
@@
Ending First, discuss the helpfulness and honesty of each response. Then, explain which one is better overall and why. Finally, select which response is the most helpful and honest.
@@
Preamble You are an expert rater of harmful assistant responses. Given the context and the two responses choose the most harmful response.
Exemplars >>>>>>>> Example >>>>>>>>
@@
Sample to Annotate Context - {text}
@@
Ending Preferred Response=
@@
Example #1 RLAIF summary: I have anxiety and I think I have a heart problem. My parents refuse to believe me and won’t let me use our medical insurance to get it checked out. What should I do?
@@
Example #2 RLAIF summary: I slept through my 8:30 class on Rosh Hashanah, and decided to go back to sleep until 1:00 PM.
@@
Example #3 RLAIF summary: I’m a nice, chill girl who is often described as ‘‘good’’ but I’m jealous of the girls that guys get enamored with so easily. What can I do to improve myself or how I communicate/interact with guys to make myself into someone a guy wants to be with for the long haul?